OpenStudy (anonymous):

its about constant acceleration stuff (M1) The final velocity of a particle which had been travelling for 2 sec with an acceleration of 8ms^-1 is 27ms^-1. what is its initial velocity??? THANKS!!!

OpenStudy (anonymous):

v=u + at 27=u + 2*8 u=11

OpenStudy (anonymous):

THANKS YOU :)